Word problem.
Ships. After a small oil spill, two cleanup ships are sent to siphon off the oil floating in Baffin Bay. The newer ship can clean up the entire spill by itself in 3 hours less time than the older ship takes by itself. Working together the two ships can clean up the oil spill in 8 hours. How long will it take the newer ship by itself to clean up the spill? Please explain how you get the answer. Thanks

Answer by stanbon(75887)

The newer ship can clean up the entire spill by itself in 3 hours less time than the older ship takes by itself. Working together the two ships can clean up the oil spill in 8 hours. How long will it take the newer ship by itself to clean up the spill?
---
older ship time = x hrs/job ; rate = 1/x job/hr
---
newer ship time = x-3 hrs/job ; rate = 1/(x-3) job/hr
------------
Together time = 8 hrs/job ; rate = 1/8 job/hr
=============================
Equation:
rate + rate = together rate
1/x + 1/(x-3) = 1/8
---
Multiply thru by 8x(x-3) to get:
8(x-3) + 8x = x(x-3)
16x-24 = x^2-3x
x^2-19x+24 = 0
--
x = [19 +- sqrt(19^2-4*24)]/2
---
x = [19 +- sqrt(265)]/2
x = [19 +- 16.28]/2
Reasonable solution:
x = 17.64 hrs (older ship time)
x-3 = 14.64 hrs (newer ship time)
============
Cheers,
Stan H.
================